OpenStudy (anonymous):

Help with: (sec^2x)(csc^2x) - sec^2x - csc^2x + 1?

OpenStudy (anonymous):

I'm not really sure how you can factor this.

OpenStudy (hhmorris98):

sec(2x)csc(2x)−sec(2x)−csc(2x+1)

Nnesha (nnesha):

rewrite all sec and csc in terms of cos and sin

Nnesha (nnesha):

what's the reciprocal of sec ?

OpenStudy (anonymous):

1/cos

Nnesha (nnesha):

correct so since it's sec^2x it would be 1/cos^2x what aabout csc^2x ?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sin^2 x }\]

OpenStudy (anonymous):

So: \[\frac{ 1 }{ \cos^2 x } x \frac{ 1 }{ \sin^2 x } - \frac{ 1 }{ \cos^2 x } - \frac{ 1 }{ \sin^2 x }\]

Nnesha (nnesha):

correct \[\rm \frac{ 1 }{ \cos^2x}*\frac{1}{\sin^2x}-\frac{1}{\cos^2x}-\frac{1}{\sin^2x}+1\] now find the common denominator there isn't any common factor

OpenStudy (anonymous):

LCD, so \[\frac{ 1 }{ \cos^2 x \sin^2 x }\]

OpenStudy (anonymous):

Actually numerator wouldn't be one.

Nnesha (nnesha):

right denominator is correct remember when the denominators are different you should multiply the top and bottom to make the same denominator

Nnesha (nnesha):

\[\rm \frac{ 1 }{ \cos^2x*sin^2x}-\frac{1}{\cos^2x}-\frac{1}{\sin^2x}+1\] to get the same denominator for 2nd fraction what would you multiply by ?

OpenStudy (anonymous):

sin^2x

Nnesha (nnesha):

correct multiply 1/cos^2x by sin^2/sin^2x \[\rm \frac{ 1 }{ \cos^2x*sin^2x}-\frac{1*\color{Red}{sin^2x}}{\cos^2x*\color{red}{sin^2x}}-\frac{1}{\sin^2x}+1\]

OpenStudy (anonymous):

So: \[\frac{ 1 }{ \cos^2 x \sin^2x } - \frac{ \sin^2x }{ \cos^2x \sin^2x } - \frac{ \cos^2x }{ \cos^2x \sin^2x } + \frac{ \cos^2x \sin^2x }{ \cos^2x \sin^2x }\] I assume you can turn the one into a fraction to make it easier?

Nnesha (nnesha):

correct and yes now write it as a single fraction :=))

OpenStudy (anonymous):

\[\frac{ 1 - \sin^2x - \cos^2x (\cos^2x) (\sin^2x) }{ \cos^2x \sin^2x }\]

OpenStudy (anonymous):

Oops, plus.

Nnesha (nnesha):

\[\frac{ 1 - \sin^2x - \cos^2x +(\cos^2x) (\sin^2x) }{ \cos^2x \sin^2x }\] like this got you

Nnesha (nnesha):

alright now use the identity sin^2x+cos^2x=1 solve for either sin or cos doesn't matter and then substitute that

OpenStudy (anonymous):

Wait, is it just going to end up being 1?

Nnesha (nnesha):

maybe how did you get one ? ;)

OpenStudy (anonymous):

Crossing out: the cos^2xsin^2x on the numerator and denominator cross out. Leaving: \[1 - \sin^2x - \cos^2x\]

Nnesha (nnesha):

you're on a right track but no you can't cancel out the sin^2 cos^2x first bec 1-sin^2x-cos^2 is also divided by cos^2sin2

Nnesha (nnesha):

first solve this identity for sin or cos sin^2x+cos^2x=1 so sin^2x= ??

OpenStudy (anonymous):

1 - cos^2x

Nnesha (nnesha):

right we can replace sin^2x with 1-cos^2x \[\frac{ 1 - (1-sin^2x)- \cos^2x +(\cos^2x) (\sin^2x) }{ \cos^2x \sin^2x }\] distribute by negative sign and then simplify

OpenStudy (anonymous):

Ugh, sorry my internet went out. I'm back.

Nnesha (nnesha):

its okay you fine :=)

OpenStudy (anonymous):

\[\frac{ -1 + \sin^2x - \cos^2x + (\cos^2x) (\sin^2x) }{ \cos^2x \sin^2x }\]

Nnesha (nnesha):

ohh wait idid a mistake it supposed to be 1-cos^2

Nnesha (nnesha):

\[\rm \frac{ 1 - (\color{Red}{1-cos^2x})- \cos^2x +(\cos^2x) (\sin^2x) }{ \cos^2x \sin^2x }\] now distribute -(1-cos^2x)

Nnesha (nnesha):

it is same as -1(1-cos^2x)

OpenStudy (anonymous):

-1 + cos^2x

Nnesha (nnesha):

right \[\rm \frac{ 1 - \color{Red}{1+cos^2x}- \cos^2x +(\cos^2x) (\sin^2x) }{ \cos^2x \sin^2x }\] and now simplify(combine like terms )

OpenStudy (anonymous):

\[\frac{ 1 + \cos^2x \sin^2x }{ \cos^2x \sin^2x }\] Then they finally cross out?

Nnesha (nnesha):

well 1 is also cancels out

Nnesha (nnesha):

\[\rm \frac{ \cancel{1} \cancel{- \color{Red}{1}+\cancel{cos^2x}}- \cancel{\cos^2x} +(\cos^2x) (\sin^2x) }{ \cos^2x \sin^2x }\]

Nnesha (nnesha):

then u left with sin^2cos^2x/sin^2cos^2x

OpenStudy (anonymous):

Okay! I get it now, thank you!

Nnesha (nnesha):

np yw :=)) good work!

Nnesha (nnesha):

well one of those lel go with np :P

OpenStudy (anonymous):

Haha.